OpenStudy (anonymous):

PLEASE HELP WITH THIS PRE-CAL QUESTION... The equation of line n is 8x-14y+3=0 What is k if the graphs of n and kx-7y+10=0 are perpendicular?

OpenStudy (anonymous):

Have you gotten any progress with this or are you just stumped?

OpenStudy (anonymous):

\[8x-14y+3=0\] I turned it into slope intertcept form \[y=\frac{ 4 }{ 7 } x+\frac{ 3 }{ 14 }\] I don't exactly know what to do now

OpenStudy (anonymous):

change the other equation into slope intercept form

OpenStudy (anonymous):

Do you know the relationship between slopes that are perpendicular?

OpenStudy (anonymous):

perpendicular means that it has the exact opposite slope y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

I got y = k/7 x + 10/7

OpenStudy (anonymous):

@killerdime20 Not exactly.

OpenStudy (anonymous):

i know that it has something to do with the slope and its perpendicular is using the reversed and using the opposite sign. Right?

OpenStudy (anonymous):

The negative reciprocal, yes. That means that if you set the negative reciprocal of the slope in y=4x/7+3/14 and set it equal to k/7, you can solve for k.

OpenStudy (anonymous):

thats what i meant. sorry for not being clear about it

OpenStudy (anonymous):

I'm wondering where the k/7 came from though

OpenStudy (anonymous):

k/7 is the slope of the perpendicular equation that you derived.

OpenStudy (anonymous):

Oh yeah I understand now

OpenStudy (anonymous):

Theoretically, if you find the negative reciprocal of A's slope, you'll have the slope of B.

OpenStudy (anonymous):

So Do I set them equal to eachother?

OpenStudy (anonymous):

It should look like this: